there is a rectangle (quadrilateral) whose width is 12cm and there is no height given..but inside the rectangle there is a diagonal line making the quadrilateral 2 equal angle.and the said diagonal line is 16cm,what is the perimeter of quadrilateral?
Answer by rothauserc(4718)

a rectangle has a length(l) and width(w), therefore its
Perimeter(P) = 2l + 2w
we are given, w = 12 and the rectangle's diagonal is 16 cm
we can use the Pythagorean Theorem to find l
l^2 + 12^2 = 16^2
l^2 + 144 = 256
l^2 = 256 -144 = 112
l = 10.583005244
P = (2*10.583005244) + (2*12) = 45.166010488 cm